OpenStudy (theedwardsfamily):

The total snowfall in Burke, Virginia, this January was 7 1/4 in. In February, 20% more snow fell than in January. How much snow fell in February in Burke, Virginia? Express your answer as a mixed number in simplest form.

OpenStudy (theedwardsfamily):

@bruhhh

OpenStudy (theedwardsfamily):

@bohotness

OpenStudy (anonymous):

Sorry, had to step away! First you convert the mixed number (7 1/4) into a normal fraction. Multiply 7 by the denominator, then add the 1 up top, so the fraction is 29/4 Now it's saying 20% (which can also be expressed as .20) MORE snow fell than 29/4. So it would look like 29/4+(.20 times 29/4) Make sense? :)

OpenStudy (anonymous):

So remember when multiplying an integer/decimal, it looks like \[\frac{ 29 }{ 4 }\times \frac{ .20 }{ 1 }=\frac{ 5.8 }{ 4 }\] At this point you wanna add the answer back to the original, so \[\frac{ 5.8+29 }{ 4 }=\frac{ 34.8 }{ 4 }\] And you can't forget to put it back in a mixed number :) \[\frac{ 34.8 }{ 4 }= 8\frac{ 7 }{ 10 }\] So 8 7/10 is your answer, I hope that helps!!

OpenStudy (theedwardsfamily):

thanks so much!!!!!!!!!!!!!!!!!!!!! :)